# DP-SPRT: Differentially Private Sequential Probability Ratio Tests

### v1

@@ 5 On DP-SPRT with Laplace Noise
$N(\theta_{0},\theta_{1},\beta,\gamma)$ $\leq\frac{2\log(1/(\gamma\beta))}{\mathrm{KL}(\nu_{\theta_{0}},\nu_{\theta_{1}})}+\frac{24(\theta_{1}-\theta_{0})\log(\zeta(s)/(1-\gamma)\beta)}{\varepsilon\mathrm{KL}(\nu_{\theta_{0}},\nu_{\theta_{1}})}+o_{\beta\rightarrow 0}\Big{(}\log(1/\beta)\Big{)}$
@@
Comparison with a Lower Bound. In Appendix F, we prove the following lower bound on the sample complexity of any $(\alpha,\beta)$-correct test that is $\varepsilon$-DP. This lower bound is valid for any distributions $\nu_{\theta}$, and follows from techniques used to prove lower bounds under DP constraints in the multi-armed bandit literature (e.g., Azize et al., (2023); Azize, (2024)).
@@
$\mathbb{E}_{\theta_{0}}[\tau]\geq\frac{\mathrm{kl}(\alpha,1-\beta)}{\min\left(\mathrm{KL}(\nu_{\theta_{0}},\nu_{\theta_{1}}),\varepsilon\cdot\mathrm{TV}(\nu_{\theta_{0}},\nu_{\theta_{1}})\right)}\text{ and }\ \mathbb{E}_{\theta_{1}}[\tau]\geq\frac{\mathrm{kl}(\beta,1-\alpha)}{\min\left(\mathrm{KL}(\nu_{\theta_{1}},\nu_{\theta_{0}}),\varepsilon\cdot\mathrm{TV}(\nu_{\theta_{0}},\nu_{\theta_{1}})\right)}$
@@
$\limsup_{\beta\rightarrow 0}\frac{\mathbb{E}_{\theta_{0}}\left[\tau\right]}{\log(1/\beta)}\leq 48\max\left(\frac{1}{\mathrm{KL}(\nu_{\theta_{0}},\nu_{\theta_{1}})},\frac{(\theta_{1}-\theta_{0})}{\mathrm{KL}(\nu_{\theta_{0}},\nu_{\theta_{1}})\varepsilon}\right)\,$,
@@
Privacy Amplification by Subsampling. DP-SPRT with Laplace noise can be further enhanced using subsampling (Steinke,, 2022). That is, in each round the empirical mean is computed based on a random subset of the available observations instead of all observations. This approach allows adding less noise for the same privacy guarantee, and in turn, a smaller correction factor. This is particularly effective in high privacy regimes as demonstrated in the experiments. Details of the implementation of this approach are provided in Appendix G.
@@ 6 Experimental Analysis
Experimental Setup. We conduct experiments on simulated Bernoulli data to evaluate the practical performance of our DP-SPRT methods. We implement the classical SPRT, PrivSPRT (Zhang et al.,, 2022), and three variants of our approach: DP-SPRT with Laplace noise, with Gaussian noise, and with Laplace noise and subsampling (Laplace-Sub). For all methods, we set target error rates $\alpha=\beta=0.05$ and perform 1000 independent trials to estimate the error probabilities and the sample complexity. We present results for Bernoulli distributions with parameter $p_{0}=0.3$ and $p_{1}=0.7$.
@@
Results. (1) Smaller Stopping Time across $\varepsilon$. Figure 1(a) shows that as privacy requirements become stringent (smaller $\varepsilon$), all private methods require more samples. Our DP-SPRT variants consistently outperform or match PrivSPRT at equivalent privacy levels, with subsampling providing substantial benefits in high-privacy regimes. This contradicts prior beliefs that Laplace noise is unsuitable for sequential testing (Zhang et al.,, 2022). (2) Smaller Probability of Error across $\varepsilon$. Figure 1(b) demonstrates that our DP-SPRT variants maintain error probabilities (significantly) below target across all privacy levels. In contrast, PrivSPRT occasionally exceeds the target rate, likely due to the noise introduced during its Monte Carlo calibration process. The fact that the empirical type I error of DP-SPRT is much below the target $\alpha$ suggests that our proposed correction function $C(n,\delta)$ is quite conservative. In Appendix H, we provide a more detailed experimental analysis with different $p_{0}$ and $p_{1}$, the variance of the stopping time, its dependence on the privacy parameters and the target error probabilities, and a tuning of the correction function.
@@ Appendix D Correctness of DP-SPRT for Exponential Families: Proof of Theorem 3
We prove the bound on type I error; the type II error follows analogously. Let $\gamma\in(0,1)$ be an arbitrary parameter that allocates the error probability.
@@
$\qquad\frac{-\mathrm{KL}(\nu_{\theta_{1}},\nu_{\theta_{0}})+\log(1/(\gamma\alpha))/n}{\theta_{1}-\theta_{0}}+\frac{Z}{n}+C\left(n,\frac{(1-\gamma)\alpha}{3}\right)\bigg{)}.$
@@
For the second term, using the condition on the correction function and a union bound,
@@
$\leq\sum_{n=1}^{\infty}\frac{(1-\gamma)\alpha}{n^{s}\zeta(s)}$
$=(1-\gamma)\alpha\sum_{n=1}^{\infty}\frac{1}{n^{s}\zeta(s)}$
$=(1-\gamma)\alpha.$

### v2

@@ 5 On DP-SPRT with Laplace Noise
$N(\theta_{0},\theta_{1},\beta,\gamma)$ $\leq\frac{24(\theta_{1}-\theta_{0})\log(\zeta(s)/(1-\gamma)\beta)}{\varepsilon\mathrm{KL}(\nu_{\theta_{0}},\nu_{\theta_{1}})}+\frac{2\log(1/(\gamma\beta))}{\mathrm{KL}(\nu_{\theta_{0}},\nu_{\theta_{1}})}+o_{\beta\rightarrow 0}\Big(\log(1/\beta)\Big).$
@@
5.1 Comparison with a Lower Bound
In Appendix F, we prove the following lower bound on the sample complexity of any $(\alpha,\beta)$-correct test that is $\varepsilon$-DP. This lower bound is valid for any distributions $\nu_{\theta}$, and follows from techniques used to prove lower bounds under DP constraints in the multi-armed bandit literature (e.g., Azize et al., (2023); Azize, (2024)).
@@ 5.1 Comparison with a Lower Bound
$\mathbb{E}_{\theta_{0}}[\tau]$ $\geq\frac{\mathrm{kl}(\alpha,1-\beta)}{\min\left(\mathrm{KL}(\nu_{\theta_{0}},\nu_{\theta_{1}}),\varepsilon\cdot\mathrm{TV}(\nu_{\theta_{0}},\nu_{\theta_{1}})\right)}\text{ and }$
$\mathbb{E}_{\theta_{1}}[\tau]$ $\geq\frac{\mathrm{kl}(\beta,1-\alpha)}{\min\left(\mathrm{KL}(\nu_{\theta_{1}},\nu_{\theta_{0}}),\varepsilon\cdot\mathrm{TV}(\nu_{\theta_{0}},\nu_{\theta_{1}})\right)}$
@@
$\limsup_{\beta\rightarrow 0}\frac{\mathbb{E}_{\theta_{0}}\left[\tau\right]}{\log(1/\beta)}\leq\frac{48}{\mathrm{KL}(\nu_{\theta_{0}},\nu_{\theta_{1}})}\max\left(1,\frac{\theta_{1}-\theta_{0}}{\varepsilon}\right)\,$,
@@
5.2 Privacy Amplification by Subsampling
DP-SPRT with Laplace noise can be further enhanced using subsampling (Steinke,, 2022). That is, in each round the empirical mean is computed based on a random subset of the available observations instead of all observations. This approach allows adding less noise for the same privacy guarantee, and in turn, a smaller correction factor. This is particularly effective in high privacy regimes as demonstrated in the experiments. Details of the implementation of this approach are provided in Appendix G.
@@ 6 Experimental Analysis
6.1 Experimental Setup
We conduct experiments on simulated Bernoulli data to evaluate the practical performance of our DP-SPRT methods. We implement the classical SPRT, PrivSPRT (Zhang et al.,, 2022), and three variants of our approach: DP-SPRT with Laplace noise, with Gaussian noise, and with Laplace noise and subsampling (Laplace-Sub). For all methods, we set target error rates $\alpha=\beta=0.05$ and perform 1000 independent trials to estimate the error probabilities and the sample complexity. We present results for Bernoulli distributions with parameter $p_{0}=0.3$ and $p_{1}=0.7$.
@@ 6.1 Experimental Setup
6.2 Results
(1) Smaller Stopping Time across $\varepsilon$. Figure 1(a) shows that as privacy requirements become stringent (smaller $\varepsilon$), all private methods require more samples. Our DP-SPRT variants consistently outperform or match PrivSPRT at equivalent privacy levels, with subsampling providing substantial benefits in high-privacy regimes. This contradicts prior beliefs that Laplace noise is unsuitable for sequential testing (Zhang et al.,, 2022).
(2) Smaller Probability of Error across $\varepsilon$. Figure 1(b) demonstrates that our DP-SPRT variants maintain error probabilities (significantly) below target across all privacy levels. In contrast, PrivSPRT occasionally exceeds the target rate, likely due to the noise introduced during its Monte Carlo calibration process. The fact that the empirical type I error of DP-SPRT is much below the target $\alpha$ suggests that our proposed correction function $C(n,\delta)$ is quite conservative. In Appendix H, we provide a more detailed experimental analysis with different $p_{0}$ and $p_{1}$, the variance of the stopping time, its dependence on the privacy parameters and the target error probabilities, and a tuning of the correction function.
@@ Appendix D Correctness of DP-SPRT for Exponential Families: Proof of Theorem 3
Type I Error. We prove the bound on type I error first. Let $\gamma\in(0,1)$ be an arbitrary parameter that allocates the error probability.
@@
$\qquad\frac{-\mathrm{KL}(\nu_{\theta_{1}},\nu_{\theta_{0}})+\log(1/(\gamma\alpha))/n}{\theta_{1}-\theta_{0}}+\frac{Z}{n}+C\left(n,(1-\gamma)\alpha\right)\bigg).$
@@
For the second term, using condition (3) and a union bound,
@@
$\leq(1-\gamma)\alpha.$
@@
Type II Error. We now prove the bound on type II error. We have
$\mathbb{P}_{\theta_{1}}(\hat{d}=0)$ $\leq\mathbb{P}_{\theta_{1}}(\tau_{0}<\infty)$
$=\mathbb{P}_{\theta_{1}}\bigg(\exists n\in\mathbb{N}_{>0}:\frac{\sum_{i=1}^{n}X_{i}}{n}-\mu_{0}+\frac{Y_{n}}{n}\leq$
$\qquad\qquad\frac{\mathrm{KL}(\nu_{\theta_{0}},\nu_{\theta_{1}})-\log(1/(\gamma\beta))/n}{\theta_{1}-\theta_{0}}-\frac{Z}{n}-C(n,(1-\gamma)\beta)\bigg)$
By the union bound, we can separate this into two events:
$\mathbb{P}_{\theta_{1}}(\hat{d}=0)$ $\leq\mathbb{P}_{\theta_{1}}\bigg(\exists n\in\mathbb{N}_{>0}:\frac{\sum_{i=1}^{n}X_{i}}{n}-\mu_{0}\leq\frac{\mathrm{KL}(\nu_{\theta_{0}},\nu_{\theta_{1}})-\log(1/(\gamma\beta))/n}{\theta_{1}-\theta_{0}}\bigg)$
$\quad+\mathbb{P}_{\theta_{1}}\bigg(\exists n\in\mathbb{N}_{>0}:\frac{Y_{n}}{n}+\frac{Z}{n}\leq-C(n,(1-\gamma)\beta)\bigg)$
For the first term, we recognize the probability of eventually accepting $H_{0}$ in the non-private SPRT when the truth is $H_{1}$. By a similar change-of-measure argument as for type I error, this probability is bounded by $\gamma\beta$.
For the second term, using condition (4), we have:
$\mathbb{P}_{\theta_{1}}\bigg(\exists n\in\mathbb{N}_{>0}:\frac{Y_{n}}{n}+\frac{Z}{n}\leq-C(n,(1-\gamma)\beta)\bigg)$ $\leq\sum_{n=1}^{\infty}\mathbb{P}_{\theta_{1}}\bigg(\frac{Y_{n}}{n}+\frac{Z}{n}<-C(n,(1-\gamma)\beta)\bigg)$
$\leq(1-\gamma)\beta$
Therefore $\mathbb{P}_{\theta_{1}}(\hat{d}=0)\leq\gamma\beta+(1-\gamma)\beta=\beta$.
Finally, we note that if the distribution $\mathcal{D}_{Y}$ is symmetric around zero and condition (3) holds, then condition (4) is automatically satisfied. This is because:
$\sum_{n=1}^{\infty}\mathbb{P}\left(\frac{Y_{n}}{n}+\frac{Z}{n}<-C(n,\delta)\right)$ $=\sum_{n=1}^{\infty}\mathbb{P}\left(-\frac{Y_{n}}{n}-\frac{Z}{n}>C(n,\delta)\right)$
$=\sum_{n=1}^{\infty}\mathbb{P}\left(\frac{Y_{n}}{n}-\frac{Z}{n}>C(n,\delta)\right)\leq\delta$
where the second equality uses the symmetry of $Y_{n}$ around zero. ∎